The joint density function of two continuous random variable X and Y is

f(x, y) = (cx2 + xy/3 , : if 0 ⩽ x ⩽ 1, 0 ⩽ y ⩽ 2

( 0, : otherwise

(a) Find c.

(b) Find the marginal pdf of X and Y .

(c) Are X and Y independent (justify)? 


1
Expert's answer
2021-09-19T14:38:50-0400

a.

Since it is a pdf then,

"\int"y"\int"x f(x, y)dxdy=1

"\intop"02"\int"01 (cx2+xy/3) dxdy=1

"\intop"02(( cx3/3+x2y/6)|01)dy=1

"\int"02 (c/3+y/6)dy =1

(cy/3+y2/12)|02=1

=2c/3+4/12=1

2c/3=2/3

c=1

therefore, the joint density function is,

f(x, y) = (x+ x*y/3 , : if 0 ⩽ x ⩽ 1, 0 ⩽ y ⩽ 2

( 0, : otherwise

b.

Let g(x) be the marginal distribution for random variable X, then its marginal distribution is defined by,

g(x)= "\int"02 f(x, y)dy

="\int"02 (x+ x*y/3)dy

=(x2y+xy2/6)|02

=2x2+4x/6

therefore,

g(x) = (2x2+4x/6, : if 0⩽ x ⩽ 1,

( 0, : otherwise

Let h(y) be the marginal distribution for random variable Y, then its marginal distribution is defined by,

h(y)= "\int"01 f(x, y)dx

= "\int"01 (x+ x*y/3)dx

=(x3/3+x2y/6)|01

=1/3+y/6

therefore,

h(y)= (1/3+y/6, : if 0 ⩽ y ⩽ 2

( 0, : otherwise

c.

To test for independence then the condition, E(XY)=E(X)*E(Y) must hold.

E(XY) = "\int"y "\int"x (x*y)* (f(x, y))dxdy

="\int"02 "\int"01 x*y*(x+ x*y/3)dxdy

="\int"02 "\int"01(x3y+(x*y)2/3)dxdy

="\int"02(x4y/4+x3y2/9)|01dy

="\int"02(y/4+y2/9)dy

=(y2/8+y3/27)|02

=4/8+8/27

=172/216

=43/54

E(X)= "\int"01 x*g(x)dx

="\int"01 x(2x2+4x/6)dx

="\int"01(2x3+4x2/6)dx

=(2x4/4+4x3/18)|01

=2/4+4/18

=26/36

=13/18

and

E(Y)="\int"02 y*h(y)dy

="\int"02 y(1/3+y/6)dy

="\int"02(y/3+y2/6)dy

=(y2/6+y3/18)|02

=4/6+8/18

=20/18

=10/9

Let us know verify if the condition hold,

E(XY)= 43/54 and E(X)*E(Y)=13/18*10/9=65/81

Since E(XY) is not equal to E(X)*E(Y), then random variables X and Y are not independent.